A 50g wooden is placed on a surface which is inclined 350 with the horizontal. Neglecting friction, what force will give the box an acceleration of 1.2 m/s ^ 2 up the incline plane.
1
Expert's answer
2020-12-08T10:53:39-0500

According to the Second Newton's Low


"F-mg\sin\alpha=ma"


So,


"F=ma+mg\sin\alpha=0.05\cdot1.2+0.05\cdot9.81\cdot\sin35°=0.34(N)" . Answer